OpenStudy (anonymous):

Show that it is not true that for every error 1/m there exists n such that |Sum(k=1,n,r^k-r/(1-r)|<1/m for all r in 0 < r < 1. I know that Sum(k=1,n,r^k) = r/(1-r)-r^(n+1)/(1-r), so the above reduces to r^(n+1) < (1-r)/m However, I'm not sure where to go from here. Please help!

OpenStudy (anonymous):

\[|\sum_{k=1}^nr^k-\frac{r}{1-r}|<\frac{1}{m}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

as you said this is the same as \[\frac{r^{n+1}}{1-r}<\frac{1}{m}\]

OpenStudy (anonymous):

so i think the important thing here is to know how to negate the statement "for every error 1/m there exists n such that |Sum(k=1,n,r^k-r/(1-r)|<1/m for all r in 0 < r < 1"

OpenStudy (anonymous):

Right. Would simply showing that n has a dependence on r be enough to disprove the statement?

OpenStudy (anonymous):

yes that is exactly the point because if this were true it would work for all \(r\)

OpenStudy (anonymous):

Ok. So would getting to \[n<\log_{r} (1-r)-\log_{r} m-1\] be enough?

OpenStudy (anonymous):

oh you are ahead of me. maybe you can use that

OpenStudy (anonymous):

the negation would be there exists an \(m\) such that for all \(n\) there exists an \(r\) such that \(\frac{r^{n+1}}{1-r}\geq\frac{1}{m}\)

OpenStudy (anonymous):

so if you can find an \(r\) that would satisfy that inequality,you have proved what you need

OpenStudy (anonymous):

that amounts to saying that \(r\) would depend on \(n\) but the proof you get to take any \(n\) and find some \(r\) that violates that condition

OpenStudy (anonymous):

So what I need to do is set the values for r and n and then find a value for m that makes the problem statement false?

OpenStudy (anonymous):

i would let \(m\) be given and show that for any \(n\) we can find an \(r\) that violates that condition

OpenStudy (anonymous):

but then i am shooting from the hip, let me think for a minute

OpenStudy (anonymous):

i mean the snappy way to say this is that the sup of \(\frac{r^{n+1}}{1-r}\) is infinity, but i don't think that is what you want. we have to be more careful

OpenStudy (anonymous):

What if I said that as r -> 1 \[\frac{ 1-r }{ m } \rightarrow 0\] and \[r ^{n+1} \ \rightarrow 1\] \[\forall m,n\] ?

OpenStudy (anonymous):

ok i had to think for a minute what you are being asked to show is that this series is convergent, but not uniformly convergent. i am not sure if you have got to that point yet, but no matter, we can solve this i think

OpenStudy (anonymous):

we pick an \(m\) say \(m=10\) although it hardly matters.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

we want to show that given any \(n\) we can find an \(r\) such that \(\frac{r^{n+1}}{1-r}\geq \frac{1}{10}\)

OpenStudy (anonymous):

and we pick our \(r\) depending on \(n\) so we can pick \(r=1-\frac{1}{n+1}\)

OpenStudy (anonymous):

in simple english, we are picking an \(r\) close to 1, which will make the denominator close to 0 and so the fraction \(\frac{r^{n+1}}{1-r}\) will be large, or at least larger than \(\frac{1}{10}\)

OpenStudy (anonymous):

we get \[\frac{r^{n+1}}{1-r}=(n+1)(1-\frac{1}{n+1})^{n+1}\] and that should be good enough, although it is now up to you to show that that is larger than \(\frac{1}{10}\)

OpenStudy (anonymous):

How did you get \[(n+1)(1-\frac{ 1 }{ n+1})^{n+1}\]?